SP: Aşağıdaki integrallerini hesaplayınız. z arctan(x)dx a) ſta b) [ 1 (1+x2) -da
Answer #1

.. In tenndni use. Integreel 134 parit. I I = fn tann da= tantno na 22 1 If itna dn 2 2 2²tanin n²+1-1 (1412) du n2 tann d nMow b leiven Iss du (1+02) take I put n = tant (1+01212 dne I dt lost Now. (1) +1212 {1+tant) Sin24 2 Cosat 1 Costt 1 (1402)tan (1) = 1 2 + tan (1) = 3T 위녁 + J-6 Sin 3-4+ * in 2* - J] 1 Sin 371 = i + - ) sin II 3 ㅋ 디어 8 2 3 IT 2 十T 2 + - 2. 4. 9T